OpenStudy (anonymous):

What is the slope of 3y=-7x? I got -2.333 Am I right?

OpenStudy (tylerd):

yep

OpenStudy (kropot72):

The exact slope is: \[\large -2 \frac{1}{3}\]

OpenStudy (anonymous):

\[y=\frac{ -7 }{ 3 } x\]

OpenStudy (anonymous):

Is the x-intercept = 0/7 = 0.00000?

OpenStudy (kropot72):

Correct.